OpenStudy (anonymous):

A car is moving at 33.0 mi/hr . Express its speed in m/s, f/s.

OpenStudy (anonymous):

14.75m/s, 4.5f/s

OpenStudy (anonymous):

1 miles = 1.609344 kilometers 1 hr = 3600 seconds 1 mile = 5280 feet then mi/hr into m/s: 33.0 mi/hr = (33.0 x 1.609344 x 1000)/3600s = .... m/s mi/hr into f/s: 33.0 mi/hr = (33.0 x 5280)/3600s = ..... f/s

OpenStudy (anonymous):

thanks!